I am a three digit number less than 200. When I am divided by 2,3,6,7 and 9,
there is a remainder of 1. What bumber am I?

Answer by Edwin McCravy(20086)

In order to leave a remainder of 1 when divided by any of
2,3,6,7, and 9, you have to be 1 more than a common multiple
of 2,3,6,7 and 9
2 = 2
3 = 3
6 = 2*3
7 = 7
9 = 3*3
The least common multiple of all those is 2*3*3*7 = 126
The next common multiple of all those is twice that or 252,
which is more than 200.
But you are less than 200.
So you are 1 more than 126, so you are 127.
Edwin